OpenStudy (anonymous):

The slope of the line whose equation is x + y = 6 is

OpenStudy (stefrheart):

slope would be -1 because you just have to subtract the x on both sides and you get y=-x+6

OpenStudy (coconutjj):

Use y=mx+b y=-x+6

OpenStudy (stefrheart):

do you understand?